Solve. Round results to the nearest thousandth.
2x^2+3x-7=0

Found 2 solutions by Ave, Waleak:
Answer by Ave(106)

I like the quadratic Formula for this one
x = (-3 ± sqrt(3^2-4(2)(-7))/2(2)
x = (-3 ± sqrt(65))/4
x = (-3 ± 8.062)/4
5.062/4 = 1.266
-11.062/4 = -2.766
{-2.766, 1.266}

Answer by Waleak(2)

There are no integers that when you add them make 3 and when you multiply them -7,so i'll use a formula to find out the roots and try to find the factored form.
2x^2+3x-7=0
a=2 b=3 C=-7 ac=-14
D=b^2-4ac=9+56=65
x1,x2=(-b almighty formula D)/20
x1,2= (-3 almighty formula 65)/4
x1=(-3- square root 65)/4
x2=(-3+square root 65)/4
well,i found the roots.so now,i will use this to find out the factured form
a(x-x2)(x-x2)
a=2
x1=(-3-square root 65)/4
x2=(-3+square root 65)/4
2[x-(-3+square root 65)/4][x-(-3-square root 65)/4]
This is the factured form.